# Hyperbolic Network Latent Space Model with Learnable Curvature

### v1

@@ 2.2 Model Setup
We consider the network data where the edges are binary and undirected without any self-loop. Such networks, consisting of $n$ nodes, can be represented as a binary adjacency matrix $A_{n\times n}$, where $A_{ij}=1$ if nodes $i$ and $j$ are connected or $A_{ij}=0$ otherwise. Thus $A$ is symmetric and all diagonal entries are 00.
Following the latent space model framework, each node $i$ is assigned with a latent position $z_{i}$, where the latent space is a $d$-dimensional hyperbolic space with curvature $-K<0$. The latent positions are treated as fixed parameters that need to be estimated. Nevertheless, our estimation procedure and theoretical results remain valid for the situation where the latent positions are generated from some stochastic procedure, in the sense of estimating the realized values of the latent position random variables.
@@
where $P_{ij}$ denotes the linkage probability between nodes $i$ and $j$, $d_{\mathbb{H}_{d}^{-K}}$ is given in (3), and $\sigma(x):[0,+\infty)\to[0,1]$ is a smooth, strictly decreasing link function such that $0\leq P_{ij}\leq 1$. Common choices of $\sigma(x)$ include logistic-type or exponential-type functions (Hoff et al.,, 2002; Lubold et al.,, 2023). For convenience, we denote $P_{n\times n}$ as the linkage probability matrix and $\Theta_{n\times n}$ as the distance matrix whose diagonal entries are 00 and off-diagonal entries are pairwise distances on the latent space, i.e., $\Theta_{ij}=\sigma^{-1}(P_{ij})=d_{\mathbb{H}_{d}^{-K}}(z_{i},z_{j})$. Writing the proposed model in the matrix form, we have:
$\Theta=\frac{1}{\sqrt{K}}\operatorname{acosh}(-Z\Lambda Z^{T}),\quad P=\sigma(\Theta),$ (5)
where $\frac{1}{\sqrt{K}}\operatorname{acosh}(x)$ and $\sigma(x)$ are applied to entries of $-Z\Lambda Z$ and $\Theta$, respectively. Equation 5 indicates that the pairwise distances are fully determined by the low-rank matrix $Z\Lambda Z^{T}$ under the hyperboloid model.
We demonstrate the advantage of the proposed hyperbolic model over the Euclidean model using a toy example in Figure 4. In this example, the network is a $2$-ary tree with a depth of 4. We fit two latent space models with hyperbolic and Euclidean distance metrics, respectively, with the link function being $\sigma(x)=2\text{logistic}(-x)=2/(1+e^{x})$. As shown in the figure, the hyperbolic embeddings, i.e., the latent positions learned from the hyperbolic model, accurately capture the hierarchical structure of the original network. In contrast, the Euclidean embeddings result in overlapping positions for four blue nodes with the yellow nodes. We further illustrate the heat maps of distance matrices learned from the two models and compare them to the shortest path distance of the original network. It is observed that hyperbolic embeddings recover more consistent pairwise distances than Euclidean embeddings.
Figure 4: A toy network with visualization of latent positions and heatmap of pairwise distances.
@@ 2.3 Identifiability
If the curvature parameter $K$ is fixed, the geometry of latent space is fully known. Combining 2.2 with Equation 5, we observe that $Z$ can only be identified up to an arbitrary hyperbolic rotation matrix $Q_{(d+1)\times(d+1)}$. Specifically, from $Q\Lambda Q^{T}=\Lambda$ we have $(ZQ)\Lambda(ZQ)^{T}=Z\Lambda Z^{T}$. Since equation $Q\Lambda Q^{T}=\Lambda$ has $(d+1)(d+2)/2$ effective constraints on entries of $Q$, we generally need $(d+1)d/2$ additional constraints to remove the indeterminacy of $Z$. In practice, a convenient set of identifiability constraints is to enforce $Z^{T}Z$ to be diagonal. Let $USU^{T}$ represent the eigen-decomposition of $Z\Lambda Z^{T}$, we know that $S$ contains exactly $d$ positive eigenvalues and one negative eigenvalue (Tabaghi and Dokmanić,, 2020). Thus, we can set $Z=U|S|^{1/2}\Lambda$, where the diagonal entries of $|S|^{1/2}$ are the square root of corresponding absolute values of eigenvalues in $S$.
In terms of identifying the curvature parameter $K$, 2.1 indicates that smooth distance-preserved mapping from one hyperbolic space to the other does not exist. However, this statement needs closer examination when considering a set of $n$ discrete points in the latent space. The problem of identifying $K$ is equivalent to show the non-existence of the isometric mapping $\phi:\mathbb{H}_{d}^{-K}\to\mathbb{H}_{d}^{-K^{\prime}}$ when $K\neq K^{\prime}$, with $\{z_{i}\}_{i=1}^{n}\subset\mathbb{H}_{d}^{-K}$ and $\{z^{\prime}_{i}\}_{i=1}^{n}\subset\mathbb{H}_{d}^{-K^{\prime}}$ such that $d_{\mathbb{H}_{2}^{-K}}(z_{i},z_{j})=d_{\mathbb{H}_{2}^{-K^{\prime}}}(\phi(z_{i}^{\prime}),\phi(z_{j}^{\prime}))$ holds for all $i\neq j$. Combined with Equation 3, we can rewrite the set of equations as
$\operatorname{acosh}(-z_{i}^{T}\Lambda z_{j})=d_{\mathbb{H}_{d}^{-1}}(z_{i},z_{j})=\sqrt{K/K^{\prime}}d_{\mathbb{H}_{d}^{-1}}(\phi(z_{i}^{\prime}),\phi(z_{j}^{\prime}))=\sqrt{K/K^{\prime}}\operatorname{acosh}(-z_{i}^{T}\Lambda z_{j}).$
It is challenging to rigorously characterize the solution because of the non-linearity of equations and the singularity issue of discrete geometry. In irregular cases, such as when all latent positions are identical or lie on a one-dimensional line, the curvature is not well-defined as the latent space becomes degenerated. However, intuitively when $n$ is sufficiently large we would not encounter singularity issues. Thus, the focus of our discussion is on regular cases where discrete points are appropriately distributed in the latent space.
In general, solutions exist only when the number of unknowns, $n(d+1)$, is greater than the total number of equations and constraints, $n(n+1)/2+(d+1)d/2$, where we include the identifiability constraints on $Z$ as well as the constraints of hyperboloid model. Solving this quadratic relationship, we can conclude that the equations above have solutions only when $n\leq d+1$, which indicates: (1) in general, $d+2$ discrete points are sufficient to determine the curvature of a hyperbolic space; and (2) when $K\neq K^{\prime}$, isometric mappings are possible only when the subspace which discrete points lie on degenerates.
Next, we provide an illustrative example to demonstrate that $4$ nodes are sufficient to distinguish the curvatures of different $2$-dimensional hyperbolic spaces. This example shares a similar spirit to the example presented in Robinson, (2006) for $2$-dimensional spherical spaces, which are metric spaces with constant positive curvatures. In this example, as shown in Figure 5, we consider four points $A,B,C,D\in\mathbb{H}_{2}^{-K}$ such that $A,B,C$ forms an equilateral hyperbolic triangle and $D$ is the midpoint of $AB$. Without loss of generality, we assume that $\overline{AB}=\overline{AC}=\overline{BC}=1$ and $\overline{AD}=\overline{BD}=1/2$, where $\overline{AB}$ denotes the hyperbolic distance between $A$ and $B$ on $\mathbb{H}_{2}^{K}$ and so on. Our goal is to find $A^{\prime},B^{\prime},C^{\prime},D^{\prime}\in\mathbb{H}_{2}^{-K^{\prime}}$ in another $2$-dimensional hyperbolic space such that all the pairwise distances between $A,B,C,D$ are preserved with $A^{\prime},B^{\prime},C^{\prime},D^{\prime}$. In the following arguments, we demonstrate that $A^{\prime},B^{\prime},C^{\prime},D^{\prime}$ exist only when $K^{\prime}=K$.
Figure 5: Two equilateral triangles in $\mathbb{D}_{2}^{-K}$ and $\mathbb{D}_{2}^{-K^{\prime}}$.
To start with, we construct an equilateral hyperbolic triangle $A^{\prime}B^{\prime}C^{\prime}$ in $\mathbb{H}_{2}^{-K^{\prime}}$ such that $\overline{A^{\prime}B^{\prime}}=\overline{A^{\prime}C^{\prime}}=\overline{B^{\prime}C^{\prime}}=1$. Then we let $D^{\prime}$ be the midpoint of $B^{\prime}C^{\prime}$ such that $\overline{A^{\prime}D^{\prime}}=\overline{B^{\prime}D^{\prime}}=1/2$. At this point, all pairwise distances between $A,B,C,D$ are preserved except for $\overline{CD}$. Thus, we need to examine the equation $\overline{CD}=\overline{C^{\prime}D^{\prime}}$ and with the relationship between $K$ and $K^{\prime}$. For $A,B,C,D$, we first apply the hyperbolic law of cosines (see for example, Bridson and Haefliger,, 2013) on $\Delta ACD$ and $\Delta BCD$:
$\cosh(\sqrt{K}~{}\overline{AC})=\cosh(\sqrt{K}~{}\overline{CD})\cosh(\sqrt{K}~{}\overline{AB}/2)-\sinh(\sqrt{K}~{}\overline{CD})\sinh(\sqrt{K}~{}\overline{AB}/2)\cos(\angle ADC),$
$\cosh(\sqrt{K}~{}\overline{BC})=\cosh(\sqrt{K}~{}\overline{CD})\cosh(\sqrt{K}~{}\overline{AB}/2)-\sinh(\sqrt{K}~{}\overline{CD})\sinh(\sqrt{K}~{}\overline{AB}/2)\cos(\angle BDC).$
With $\angle ADC+\angle BDC=\pi$ and $\overline{AB}=\overline{AC}=\overline{BC}=1$, by adding both sides of the two equations we have:
$\cosh(\sqrt{K})=\cosh(\sqrt{K}/2)\cosh(\sqrt{K}~{}\overline{CD}).$ (6)
Similarly, applying the hyperbolic law of cosines for $A^{\prime},B^{\prime},C^{\prime},D^{\prime}$ gives us
$\cosh(\sqrt{K^{\prime}})=\cosh(\sqrt{K^{\prime}}/2)\cosh(\sqrt{K^{\prime}}~{}\overline{C^{\prime}D^{\prime}}).$ (7)
Combining Equation 6 and Equation 7, we know that $\overline{CD}=\overline{C^{\prime}D^{\prime}}$ holds if and only if
$\frac{1}{\sqrt{K}}\cosh^{-1}\left[\frac{\cosh(\sqrt{K})}{\cosh(\sqrt{K}/2)}\right]=\frac{1}{\sqrt{K^{\prime}}}\cosh^{-1}\left[\frac{\cosh(\sqrt{K^{\prime}})}{\cosh(\sqrt{K^{\prime}}/2)}\right].$
However, function $f(K)=\frac{1}{\sqrt{K}}\cosh^{-1}\left[\frac{\cosh(\sqrt{K})}{\cosh(\sqrt{K}/2)}\right]$ is a strictly increasing function w.r.t. $K>0$. Thus, we conclude that $\overline{CD}=\overline{C^{\prime}D^{\prime}}$ holds if and only if $K=K^{\prime}$.
Our example illustrates that it is impossible to find a distance-preserved mapping for any two sets of four points from $2$-dimensional hyperbolic spaces with different curvatures. In other words, $4$ nodes are sufficient to distinguish the curvatures of different $2$-dimensional hyperbolic spaces. Moreover, analogous examples can be extended to constructions beyond equilateral triangles. Later in Section 4.1, we will directly quantify the embedding error caused by the misspecified curvature.
To summarize our main conclusions on model identifiability, we have the following observations: (1) to identify the latent position matrix $Z$, it is necessary to impose $(d+1)d/2$ constraints; and (2) to identify the curvature parameter $K$, we generally require $n\geq d+2$.
@@ 3 Estimation Procedure
We propose to estimate the model parameters by minimizing the negative log-likelihood function:
$\mathcal{L}(K,Z)=-\sum_{i\neq j}[A_{ij}\log(\sigma\left(\Theta_{ij})\right)+(1-A_{ij})\log\left(1-\sigma(\Theta_{ij})\right)],~{}\operatorname{subject~{}to}~{}z_{i}^{T}\Lambda z_{i}=-1,1\leq i\leq n,$ (8)
where $\Theta_{ij}=\operatorname{acosh}(-z_{i}^{T}\Lambda z_{j})/\sqrt{K}$. This optimization problem associated with both $K$ and $Z$ is challenging due to the non-convex function form and the constraints on $Z$. To address these challenges, we develop a gradient-based optimization algorithm outlined in Algorithm 1. In each iteration, the algorithm updates $K$ with Gradient Descent (GD) and updates $Z$ with Manifold Gradient Descent (MGD) simultaneously. Since the effective sample sizes when calculating the gradients of $K$ and $Z$ are different, we propose to use different step sizes $\eta_{K}$ and $\eta_{Z}$ in the algorithm.
@@
1:Adjacency matrix: $A$; dimension: $d$; initial values: $K^{0},Z^{0}$; step sizes: $\eta_{K},\eta_{Z}$; threshold: $\epsilon$.
@@
8:$\hat{K}=K^{t}$, $\hat{Z}=Z^{t}$.
Utilizing Riemannian optimization techniques (Nickel and Kiela,, 2018), MGD directly updates latent positions along the gradient direction on the hyperboloid manifold. Outlined in Algorithm 2, MGD first derives the direction of the steepest descent from the Euclidean gradient and then calculates the Riemannian gradient on the tangent space $\mathcal{T}_{z}=\{x\in\mathbb{R}^{d+1}|x^{T}\Lambda z=0\}$ with projection mapping (9). In the second step, MGD updates the latent position along the geodesic on $\mathbb{H}_{d}^{-K}$ with the Riemannian gradient direction, using the exponential map (10). Note that steps 4 and 5 inside the for-loop can be updated with parallel computing to accelerate computation.
$\text{proj}_{z}(g)$ $:\mathbb{R}^{d+1}\to\mathcal{T}_{z}\quad s.t.\quad\text{proj}_{z}(g)=g+(g^{T}\Lambda z)\cdot z,$ (9)
$\text{exp}_{z}(g)$ $:\mathcal{T}_{z}\to\mathbb{H}_{d}^{-K}\quad s.t.\quad\text{exp}_{z}(g)=\cosh(\sqrt{g^{T}\Lambda g})\cdot z+\sinh(\sqrt{g^{T}\Lambda g})/\sqrt{g^{T}\Lambda g}\cdot g.$ (10)
@@
1:Adjacency matrix: $A$; dimension: $d$; step size: $\eta_{Z}$; estimators: $K^{t},Z^{t}$.
@@
3:$g^{E}=\Lambda\nabla_{Z}\mathcal{L}(K^{t},L^{t};A,d)$; $\triangleright$ Calculate the steepest direction.
4:for $i=1,2,...n$ do
@@
8:$Z^{t+1}=\left(z_{i}^{t+1}\right)_{i=1}^{n}$.
Since the optimization problem is non-convex, initialization plays an important role in finding the global minimum and the corresponding optimal estimators. We use $P^{0}$, $Z^{0}$, and so on, to denote the initial values of these parameters. Following Ma et al., (2020), we develop two initialization methods based on the Universal Singular Value Thresholding (USVT, Chatterjee,, 2015). Intuitively, USVT produces a low-rank estimation $P^{0}$ of the probability matrix by taking SVD on $A$. By applying the inverse of link function $\sigma(x)$ on $P^{0}$, we are then able to derive an initial estimation $\Theta^{0}$ of the distance matrix. Nevertheless, solving both $K^{0}$ and $Z^{0}$ from $\Theta^{0}$ remains to be a challenging problem due to the non-convexity.
If $K$ is known, according to Equation 5 an initial estimation of $G_{0}=Z^{0}\Lambda{Z^{0}}^{T}$ can be obtained with an entry-wise inverse mapping. Then $Z^{0}$ can be obtained with the eigen-decomposition of $G_{0}$. Denote $G_{0}=USU^{T}$ to be the eigen-decomposition of $G_{0}$, we define $\tilde{Z}_{0}=\tilde{U}\tilde{|S|}^{1/2}\Lambda$ to be a low-rank approximation, where $\tilde{S}$ contains the $d$ leading positive eigenvalues and the smallest negative eigenvalue and $\tilde{U}$ contains the corresponding eigenvectors. Finally, the initial estimation $Z^{0}$ can be obtained by projecting $\tilde{Z}$ onto $\mathbb{H}_{d}^{-K}$ using an established formula in the literature (Tabaghi and Dokmanić,, 2020). In practice, when we do not have access to $K$, we propose to first fit the model with several initial candidate values of $K^{0}$ while fixing $\eta_{K}=0$, and choose the value with the lowest loss as $K^{0}$ and run Algorithm 1. The initialization method is summarized in Algorithm 3.
@@
1:Adjacency matrix: $A$; dimension: $d$; threshold: $\tau$; candidates list of $K^{0}$: $\{K_{1},K_{2},...,K_{m}\}$.
@@
3:$P^{0}=\sum_{\sigma_{i}\geq\tau}\sigma_{i}u_{i}v_{i}^{T}$, where $\sum_{1\leq i\leq n}\sigma_{i}u_{i}v_{i}^{T}$ is SVD of $A$; $\triangleright$ USVT
4:$\Theta^{0}=\sigma^{-1}[(P^{0}+{P^{0}}^{T})/2]$;
5:for $j=1,2,...,m$ do

### v2

@@ 2.2 Model Setup
We consider the network data where the edges are binary and undirected without any self-loop. Such networks, consisting of $n$ nodes, can be represented as a binary adjacency matrix $A_{n\times n}$, where $A_{ij}=1$ if nodes $i$ and $j$ are connected or $A_{ij}=0$ otherwise. Thus $A$ is symmetric and all diagonal entries are 0. Following the latent space model framework, each node $i$ is assigned with a latent position $z_{i}$, where the latent space is a $d$-dimensional hyperbolic space with curvature K<0-K<0. The latent positions are treated as fixed parameters that need to be estimated. Nevertheless, our estimation procedure and theoretical results remain valid for the situation where the latent positions are generated from some stochastic procedure, in the sense of estimating the realized values of the latent positions.
@@
where $P_{ij}$ denotes the linkage probability between nodes $i$ and $j$, $d_{\mathbb{H}_{d}^{-K}}$ is given in (3), and $\sigma(x):[0,+\infty)\to[0,1]$ is a smooth, strictly decreasing link function such that $0\leq P_{ij}\leq 1$. Common choices of $\sigma(x)$ include logistic-type or exponential-type functions (Hoff et al., 2002; Lubold et al., 2023). For convenience, we denote $P_{n\times n}$ as the linkage probability matrix and $\Theta_{n\times n}$ as the distance matrix whose diagonal entries are 0 and off-diagonal entries are pairwise distances on the latent space, i.e., $\Theta_{ij}=\sigma^{-1}(P_{ij})=d_{\mathbb{H}_{d}^{-K}}(z_{i},z_{j})$. Writing the proposed model in the matrix form, we have:
$\Theta=\frac{1}{\sqrt{K}}\operatorname{acosh}(-Z\Lambda Z^{T})\mbox{ and }P=\sigma(\Theta),$ (5)
where $\operatorname{acosh}(x)/{\sqrt{K}}$ and $\sigma(x)$ are applied to entries of $Z\Lambda Z$ and $\Theta$, respectively. Equation 5 indicates that the pairwise distances are fully determined by the low-rank matrix $Z\Lambda Z^{T}$ under the hyperboloid model. Note that from Equation 5, $K$ is directly related to network sparsity and various other empirical properties. It serves as a modeling parameter for network sparsity and heterogeneity, which will be further discussed in Section 5.1.
We demonstrate the advantage of the proposed hyperbolic model over the Euclidean model using a toy example in Figure 4. In this example, the network is a $2$-ary tree with a depth of 4. We fit two latent space models with hyperbolic and Euclidean distance metrics, respectively, with the link function being $\sigma(x)=2\text{logistic}(-x)=2/(1+e^{x})$. As shown in the figure, the hyperbolic embeddings, i.e., the latent positions learned from the hyperbolic model, accurately capture the hierarchical structure of the original network. In contrast, the Euclidean embeddings result in overlapping positions for four blue nodes with the yellow nodes. We further illustrate the heat maps of distance matrices learned from the two models and compare them to the shortest path distance of the original network. It is observed that hyperbolic embeddings recover more consistent pairwise distances than Euclidean embeddings. In the Supplementary Materials (Section B.3), we further provide examples demonstrating the advantage of hyperbolic models for embedding networks with community structures.
Figure 4: A toy network with visualization of latent positions and distance heatmap.
@@ 2.3 Identifiability
If the curvature parameter $K$ is fixed, the geometry of latent space is fully known. Combining 2.2 with Equation 5, we observe that $Z$ can only be identified up to an arbitrary hyperbolic rotation matrix $Q_{(d+1)\times(d+1)}$. Specifically, from $Q\Lambda Q^{T}=\Lambda$ we have $(ZQ)\Lambda(ZQ)^{T}=Z\Lambda Z^{T}$. Since equation $Q\Lambda Q^{T}=\Lambda$ has $(d+1)(d+2)/2$ effective constraints on entries of $Q$, we generally need $(d+1)d/2$ additional constraints to remove the indeterminacy of $Z$. In practice, a convenient set of identifiability constraints is to enforce $Z^{T}Z$ to be diagonal. Let $USU^{T}$ represent the eigen-decomposition of $Z\Lambda Z^{T}$, we know that $S$ contains exactly $d$ positive eigenvalues and one negative eigenvalue (Tabaghi and Dokmanić, 2020). Thus, we can set $Z=U|S|^{1/2}\Lambda$, where the diagonal entries of $|S|^{1/2}$ are the square root of corresponding absolute values of eigenvalues in $S$.
In terms of identifying the curvature parameter $K$, 2.1 indicates that smooth distance-preserved mapping from one hyperbolic space to the other does not exist. However, this statement needs closer examination when considering a set of $n$ discrete points in the latent space. The problem of identifying $K$ is equivalent to show the non-existence of the isometric mapping $\phi:\{z_{i}\}_{i=1}^{n}\subset\mathbb{H}_{d}^{-K}\to\{z^{\prime}_{i}\}_{i=1}^{n}\subset\mathbb{H}_{d}^{-K^{\prime}}$, such that $d_{\mathbb{H}_{2}^{-K}}(z_{i},z_{j})=d_{\mathbb{H}_{2}^{-K^{\prime}}}(\phi(z_{i}),\phi(z_{j}))$ holds for all $i\neq j$ when $K\neq K^{\prime}$. Combined with Equation 3, we can rewrite the set of equations as $\operatorname{acosh}(-z_{i}^{T}\Lambda z_{j})=d_{\mathbb{H}_{d}^{-1}}(z_{i},z_{j})=\sqrt{K/K^{\prime}}d_{\mathbb{H}_{d}^{-1}}(\phi(z_{i}),\phi(z_{j}))=\sqrt{K/K^{\prime}}\operatorname{acosh}(-{z_{i}^{\prime}}^{T}\Lambda{z_{j}^{\prime}}).$ It is challenging to rigorously characterize the solution because of the non-linearity of equations and the singularity issue of discrete geometry. In irregular cases, such as when all latent positions are identical or lie on a one-dimensional line, the curvature is not well-defined as the latent space becomes degenerated. However, intuitively when $n$ is sufficiently large we would not encounter singularity issues. Thus, the focus of our discussion is on regular cases where discrete points are appropriately distributed in the latent space. In general, solutions exist only when the number of unknowns, $n(d+1)$, is greater than the total number of equations and constraints, $n(n+1)/2+(d+1)d/2$, where we include the identifiability constraints on $Z$ as well as the constraints of hyperboloid model. Solving this quadratic relationship, we can conclude that the equations above have solutions only when $n\leq d+1$, which indicates: (1) in general, $d+2$ discrete points would be sufficient to determine the curvature of a hyperbolic space; and (2) when $K\neq K^{\prime}$, isometric mappings are possible only when the subspace which discrete points lie on degenerates. To further illustrate this, we provide an illustrative example to demonstrate that 4 nodes are sufficient to distinguish the curvatures of different 2-dimensional hyperbolic spaces in the Supplementary Materials.
@@ 3 Estimation Procedure
We estimate the model parameters by minimizing the negative log-likelihood function:
$\mathcal{L}(K,Z)=-\sum_{i\neq j}[A_{ij}\log(\sigma\left(\Theta_{ij})\right)+(1-A_{ij})\log\left(1-\sigma(\Theta_{ij})\right)],~\mbox{s.t.}~z_{i}^{T}\Lambda z_{i}=-1,1\leq i\leq n,$
where $\Theta_{ij}=\operatorname{acosh}(-z_{i}^{T}\Lambda z_{j})/\sqrt{K}$. This optimization problem associated with both $K$ and $Z$ is challenging due to the non-convex function form and the constraints on $Z$. To address these challenges, we develop a gradient-based optimization algorithm outlined in Algorithm 1. In each iteration, the algorithm updates $K$ with Gradient Descent (GD) and updates $Z$ with Manifold Gradient Descent (MGD) simultaneously. We propose using different step sizes for updating $K$ and $Z$ to ensure stability, with each step size normalized by the effective sample sizes when calculating the corresponding gradients. In particular, we choose $\eta_{K}=O(K^{0}/n^{2})$ and $\eta_{Z}=O(1/n)$, respectively. Notably, our numerical results show that this choice of $\eta_{K}$ also prevents the updated $K$ from being negative and invalid in Algorithm 1.
@@
1:Adjacency matrix: AA; dimension: $d$; initial values: $K^{0},Z^{0}$; step sizes: $\eta_{K},\eta_{Z}$; threshold: $\epsilon$.
@@
[…]
Utilizing Riemannian optimization techniques (Nickel and Kiela, 2018), MGD directly updates latent positions along the gradient direction on the hyperboloid manifold. Outlined in Algorithm 2, MGD first derives the direction of the steepest descent from the Euclidean gradient and then calculates the Riemannian gradient on the tangent space $\mathcal{T}_{z}=\{x\in\mathbb{R}^{d+1}|x^{T}\Lambda z=0\}$ with projection mapping (6). In the second step, MGD updates the latent position along the geodesic on $\mathbb{H}_{d}^{-K}$ with the Riemannian gradient direction, using the exponential map (7). Note that steps 4 and 5 inside the for-loop can be updated with parallel computing to accelerate computation.
$\text{proj}_{z}(g)$ $:\mathbb{R}^{d+1}\to\mathcal{T}_{z}\quad\mbox{s.t.}\quad\text{proj}_{z}(g)=g+(g^{T}\Lambda z)\cdot z,$ (6)
$\text{exp}_{z}(g)$ $:\mathcal{T}_{z}\to\mathbb{H}_{d}^{-K}\quad\mbox{s.t.}\quad\text{exp}_{z}(g)=\cosh(\sqrt{g^{T}\Lambda g})\cdot z+\sinh(\sqrt{g^{T}\Lambda g})/\sqrt{g^{T}\Lambda g}\cdot g.$ (7)
@@
1:Adjacency matrix: AA; dimension: $d$; step size: $\eta_{Z}$; estimators: $K^{t},Z^{t}$.
@@
[…]
4:for i=1,2,ni=1,2,...n do
@@
[…]
Since the optimization problem is non-convex, initialization plays an important role in finding the global minimum and the corresponding optimal estimators. We use $P^{0}$, $Z^{0}$, and so on, to denote the initial values of these parameters. Following Ma et al. (2020), we develop two initialization methods based on the Universal Singular Value Thresholding (USVT, Chatterjee, 2015). Intuitively, USVT produces a low-rank estimation $P^{0}$ of the probability matrix by taking SVD on $A$. By applying the inverse of link function $\sigma(x)$ on $P^{0}$, we are then able to derive an initial estimation $\Theta^{0}$ of the distance matrix. Nevertheless, solving both $K^{0}$ and $Z^{0}$ from $\Theta^{0}$ remains to be a challenging problem due to the non-convexity. If $K$ is known, according to Equation 5 an initial estimation of $G_{0}=Z^{0}\Lambda{Z^{0}}^{T}$ can be obtained with an entry-wise inverse mapping. Then $Z^{0}$ can be obtained with the eigen-decomposition of $G_{0}$. Denote $G_{0}=USU^{T}$ to be the eigen-decomposition of $G_{0}$, we define $\tilde{Z}_{0}=\tilde{U}\tilde{|S|}^{1/2}\Lambda$ to be a low-rank approximation, where $\tilde{S}$ contains the $d$ leading positive eigenvalues and the smallest negative eigenvalue and $\tilde{U}$ contains the corresponding eigenvectors. Finally, the initial estimation $Z^{0}$ can be obtained by projecting $\tilde{Z}$ onto $\mathbb{H}_{d}^{-K}$ using an established formula in the literature (Tabaghi and Dokmanić, 2020). In practice, when we do not have access to $K$, we propose to first fit the model with several initial candidate values of $K^{0}$ while fixing $\eta_{K}=0$, and choose the value with the lowest loss as $K^{0}$ and run Algorithm 1. The initialization method is summarized in Algorithm 3.
@@
1:Adjacency matrix: AA; dimension: $d$; threshold: $\tau$; candidates list of $K^{0}$: $\{K_{1},K_{2},...,K_{m}\}$.
@@
[…]
5:for j=1,2,,mj=1,2,...,m do